OpenStudy (anonymous):

What is the sum of the geometric sequence −4, 24, −144, ... if there are 8 terms

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

Eighty-Nineeeeeee

OpenStudy (anonymous):

−959,781 29,637 26,661 959,780

OpenStudy (anonymous):

anybody know the actual equation to find these?

OpenStudy (kirbykirby):

\[\Large \sum_{k=1}^nar^{k-1}=a\left( \frac{1-r^n}{1-r}\right) \] where \(r\) is the common ratio, which you can find by dividing the 2nd term with the 1st, or the 3rd term with the 2nd, etc.

OpenStudy (anonymous):

ohh yeah my calculator has the equation input for that, what does n stand for? is that the first term or the one we are trying to find?

OpenStudy (kirbykirby):

Well notice that \[\Large \sum_{k=1}^nar^{k-1}=a+ar+ar^2+\ldots+ar^{n-1}\\ \Large \Large \sum_{k=1}^{8}ar^{k-1}=-4+24-144+...+ar^{8-1}\]

OpenStudy (anonymous):

ok so are n is the term we are trying to find, now why does our variable = 1?

OpenStudy (kirbykirby):

well it starts at k=1, and goes until k=8, for 8 terms. You should see though from the way I wrote the sum: \(a = -4 \\ ar = 24 \) You can deduce \(r\) and then use the formula in my first post

OpenStudy (anonymous):

ooh ok gotcha, i was wandering because when i put it into my calc it has tobe right

OpenStudy (anonymous):

hmm normaly when the equation is already laid out it comes up right, but im getting large numbers, what would the expression be

OpenStudy (anonymous):

because i enter it in like this.. Expression,variable,low,high all are correct but the exp

OpenStudy (kirbykirby):

\[ \Large \sum_{k=1}^8-4(-6)^{k-1}=(-4)\left(\frac{1-(-6)^8}{1-(-6)}\right)\]

OpenStudy (anonymous):

ok i got the -4 in the exp correct but where did the 6 come from

OpenStudy (kirbykirby):

That's the \(r\), and its -6 a = -4 ar = 24 Divide the 2nd equation by the 1st: \[\frac{ar}{a}=\frac{24}{-4}\implies r = -6 \]

OpenStudy (anonymous):

oh thats right r = ratio

OpenStudy (kirbykirby):

yup !

OpenStudy (anonymous):

i think my calculator isnt reading it right i have never had it do this but its giving me \[\frac{ 239945 }{ 419904 }\]

OpenStudy (kirbykirby):

try like:(-4)*((1-(-6)^8)/(1+6))

OpenStudy (kirbykirby):

If you are typing in the inputs for some "geometric series" function on your calculator.. Well to be honest I am not sure what your calculator is doing, or how the inputs must be put in. I never used a graphing calculator, I only ever had a simple scientific calculator. I just rely on the formulas such as the one abov

OpenStudy (anonymous):

im using the TI nSpire cx cas and all i do is go to the "Glossary" and find that backwards e thing and it says to input it as such Exp.Var,Low,High

OpenStudy (anonymous):

normally i get problems that already have it written out and i just input the values to solve

OpenStudy (anonymous):

ok using ur formula i got D

OpenStudy (kirbykirby):

Hm. Well I think this is a case of "blindly" using the calculator without knowing exactly what it's doing :) Um, if you have your calculator's instructions, you should be able to see what he calculator is doing in that function, and type in the appropriate inputs accordingly.

OpenStudy (kirbykirby):

Again, I don't have this calculator so I can't help too much in regards to using it :(

OpenStudy (anonymous):

it has it written out in its description in the glossary. its almost like its not reading it for example normally when i come across these problems it looks like this \[\sum_{i=7}^{21}\](4-2)^i-1 now that example probably doesnt equal anything or even work but then i have four choices, i enter it into my calc as the example is written and it gives me the answer, i have never seen it do this before

OpenStudy (anonymous):

almost like we arent writing it like it would read it.. im baffled,, but i wish i could figure out whats wrong because i have 3 more of these problems that i need to do and i dont want to keep asking help on the same types of problems

OpenStudy (kirbykirby):

Hm give me a sec I will try and search for the instructions online for your calculator. (I do have to do you are lucky you can use it for your problems haha. But it's always good to know how to go back to basics :) )

OpenStudy (kirbykirby):

Ok so it looks like this? https://epsstore.ti.com/OA_HTML/csksxvm.jsp?nSetId=109208

OpenStudy (anonymous):

yeah but u know whats funny i did not know you could fill out the values like that ive always done it like \[\Sigma(Expression,Variable,Low,High)\]

OpenStudy (anonymous):

u showed me sumthing new about my own calc lol

OpenStudy (kirbykirby):

oh interesting haha

OpenStudy (anonymous):

it came up with that answer first try!

OpenStudy (anonymous):

when i did it like that page did it

OpenStudy (anonymous):

did not even know sigma was calc. im retaking alg 2

OpenStudy (kirbykirby):

yay :)

OpenStudy (anonymous):

but now how again did you determine six? in the expression?

OpenStudy (kirbykirby):

r is the common ratio, to do that quickly: Divide the 2nd term by the 1st

OpenStudy (kirbykirby):

so 24/(-4) = -6

OpenStudy (anonymous):

now deos dividing the second by the first always work?

OpenStudy (kirbykirby):

yes always :)

OpenStudy (anonymous):

ok im gonna work another one out and then tag u to make sure i did it correctly

OpenStudy (kirbykirby):

coolies ;)